cpl106:

Sam purchased a new car for $17,930. The value of the car depreciated by 19% per year. When he trades the car in after x years, the car is worth no more than $1,900. Fill in the values of a, b, and c to complete the exponential inequality of the form a(b)x ≤ c that can be used to determine the number of years after which the car is worth no more than $1,900.

cpl106:

cpl106:

@ramen

supie:

\[a(b)^x=c\] new car-\[$17,930\] \[ a = 17930\] value of the car is decreased by 19%/year \[ b = 1-0.19 = 0.81\] The car was valued more ≥ 1,900 \[ c = 1900\] the exponential inequality is \[ a (b) ^ x ≤ c\] \[ 17930 (0.81) ^ x ≤ 1900\]

cpl106:

thank you

supie:

np